prove divD= volume charge density using gauss law .Also find the capacitance of two concentric spherical metal shells radii a and b.

1
Expert's answer
2021-12-17T11:50:07-0500

"\oint_S\vec D dS=\iiint_V\text{div}\vec Dd V,"

"Q=\int_V\rho dV,"

"\smallint_V\text{div}\vec DdV=\int_V\rho dV,\implies"

"\text{div}\vec D=\rho."


"U=U_2-U_1=kQ(\frac 1{R_1}-\frac 1{R_2}),"

"C=\frac QU=\frac{R_1R_2}{k(R_2-R_1)}."